OpenStudy (anonymous):

@ash2326 5.Let f(x) = x4 – 8x3 + 16x2 – 19 and g(x) = x – 5. Find

OpenStudy (ash2326):

Just 5 minutes

OpenStudy (anonymous):

ok :)

OpenStudy (ash2326):

Sorry taking me long. I'll be on it ASAP

OpenStudy (anonymous):

dont worry about it take your time :)

OpenStudy (ash2326):

Hey ricky

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

ash i will get back online in 20 minutes i have to go for a little bit :/

OpenStudy (ash2326):

We have \[f(x)=x^4-8x^3+16x^2-19\] and \[g(x)=x-5\] We want to find \[\frac{f(x)}{g(x)}\] Let's divide __x^3____________________________________ x-5 | x^4-8x^3+16x^2-19 x^4-5x^3 __x^3___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3 __x^3___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 __x^3-3x^2___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -3x^3+15x^2 __x^3-3x^2___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -( -3x^3+15x^2) __x^3-3x^2___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -( -3x^3+15x^2) x^2-19 __x^3-3x^2+x___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -( -3x^3+15x^2) x^2-19 -( x^2-5x) __x^3-3x^2+x___________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -( -3x^3+15x^2) x^2-19 -(x^2-5x) 5x-19 __x^3-3x^2+x_+5_________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -( -3x^3+15x^2) x^2-19 -(x^2-5x) 5x-19 (5x-25) __x^3-3x^2+x+5_________________________________ x-5 | x^4-8x^3+16x^2-19 -( x^4-5x^3) 0 -3x^3+16x^2-19 -( -3x^3+15x^2) x^2-19 -(x^2-5x) 5x-19 - (5x-25) 6 so we get quotient \[x^3-3x^2+x+5\] and remainder 6